OpenStudy (anonymous):

Solve using the multiplication principle. -4y<=1/2

OpenStudy (anonymous):

divide both sides by \(-4\) and change \(\leq \) to \(\geq \) since \(-4\) is negative

OpenStudy (anonymous):

Hi there, I believe the answer to your problem is... \[y \ge \frac{ 1 }{ 8 }\] \(\color{red}{Chance}\)

OpenStudy (anonymous):

Can you best responce please?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."